Based on data from the Statistical Abstract of the United States, 112th Edition, only about 10% of senior citizens (65 years old or older) get the flu each year. However, about 30% of the people under 65 years old get the flu each year. In the general population, there are 15% senior citizens (65 years old or older).

(a) What is the probability that a person selected at random from the general population is senior citizen who will get the flu this season? (Use 3 decimal places.)


(b) What is the probability that a person selected at random from the general population is a person under age 65 who will get the flu this year? (Use 3 decimal places.)


(c) Answer parts (a) and (b) for a community that has 94% senior citizens. (Use 3 decimal places.)


(d) Answer parts (a) and (b) for a community that has 49% senior citizens. (Use 3 decimal places.)
Answer #1

Answer:

Given that:

Based on data from the Statistical Abstract of the United States, 112th Edition, only about 10% of senior citizens (65 years old or older) get the flu each year. However, about 30% of the people under 65 years old get the flu each year. In the general population, there are 15% senior citizens (65 years old or older).

Consider the below information

P(Senior Citizens) = 15%

= 0.15


P(Under 65) = 85

= 0.85

P(Flu /Senior Citizen) = 10%

= 0.10

P(Flu /under 65) = 30%

= 0.30

If A and B are any two events,

P(A and B) = P(A)*P(B, given A)

a) What is the probability that a person selected at random from the general population is senior citizen who will get the flu this season?

The probability  that a person selected at random from the general population is a senior citizen who will get the flu this year is,

P( A person is selected at random from the general population is a senior citizen who will get flu this year)

= P(senior citizens) .P(Flu /senior citizen)

= (0.15)*(0.10)

= 0.015


Therefore, the probability that  person selected at random from the general population is senior citizen who will get the flu this year is 0.015

b) What is the probability that a person selected at random from the general population is a person under age 65 who will get the flu this year?

The probability  that a person selected at random from the general population is a person under age 65 who will get the flu this year is,

P( A  person selected at random from the general population is a person under age 65 who will get the flu this year )

= P(Under age 65) P(flu/ Under age 65)

= (0.85)*(0.30)

= 0.255


Therefore the probability that a person selected at random from the general population is a person under age 65 who will get the this year is 0.255

c)  Answer parts (a) and (b) for a community that has 94% senior citizens

Percentage  of senior citizen is 94%.

That is, P(Senior Citizens ) = 95%

= 0.95


The probability that a person selected at random from the general population is a senior citizen who will get the flu this year is,

P( A person selected at random from the general population is a senior citizen who will get the flu this year )

= P(Senior citizens) .P(Flu/Senior Citizen)

= (0.94)*(0.10)

= 0.094


Therefore, the probability that a person selected at random from the general population senior citizen who will get theft this year is 0.094

The probability a person selected at random from the general population is a person under age 65 who will get the flu this year is,

P( A  person selected at random from the general population is a person under age 65 who will get the flu this year)

= P(Under age 65).P(Flu/Under age 65)

= ( 0.06)*(0.30)

= 0.018


Therefore, the probability that a person selected at random from the general population is a person under age 65 who will get the flu this year is 0.018

d) Answer parts (a) and (b) for a community that has 49% senior citizens

Percentage of senior citizen is 49%
That is

P(Senior Citizens) = 49%

= 0.49


The probability that a person selected at random from the general population is a senior citizen who will get the flu this year is,

P(A person selected at random from the general population is a senior citizen who will get the flu this year )

= P(Senior Citizens).P(Flu/Senior Citizens)

= (0.49) *(0.10)

= 0.049


Therefore, the probability that a person selected at random from the general population is a senior citizen who will get the flu this year is 0.049.

The probability a person selected at random from the general population is a person under age 65 who will get the flu this year is,

P( A person selected at random from the general population is a person under age 65 who will get the flu this year)

= P(Under age 65) .P(Flu/Under age 65)

= (0.49)*(0.30)

= 0.147


Therefore the probability that person selected at random from the general population is a person under age who get the flu this year is 0.147